OpenStudy (anonymous):

Factor completely xy-2x+4y-8

OpenStudy (anonymous):

xy-2x+4y-8 = 0 xy+4y-2x-8 = 0 (Bring the whole number amd the -ve terms to the end.) now take the common factors. y(x+4) -2(x+4) = 0 (x+4) (y-2) = 0 solved.

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

I might actually get passed this factoring so I can recieve my credits for next year :D